OpenStudy (fibonaccichick666):

So, just wondering about terminology in this question: Compute the nth taylor polynomial at 0 for the exponential function. "Exponential function"-Does that mean \(e^x\)? I've never heard that terminology before.

OpenStudy (fibonaccichick666):

@SithsAndGiggles

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

ok thanks!

OpenStudy (anonymous):

although sometimes it is written \(\exp(x)\)

OpenStudy (fibonaccichick666):

ah, ok that I have seen, but never exponential function

OpenStudy (fibonaccichick666):

quick follow up, I got \(P_n(x)=1+x+\frac{x^2}{2}+\frac{x^3}{3!}+...+\frac{x^n}{n!}\). It seems too simple, did I do something wrong?

OpenStudy (fibonaccichick666):

@satellite73, when you come back